# Effects of Collaboration on the Performance of Interactive Theme Discovery Systems

### v4

@@ Appendix A Topic Model Experimental Settings
To ensure comparable results, we use the same starting distribution of 13 topics for all our experiments using the same hyperparameters as fang-etal-2021-query, which are $\alpha=1.0$, $\beta=0.5$, $\gamma=1.5$. The QDTM also allows queries to be input prior to topic model initialization to produce partitions that follow prior knowledge, but we do not take advantage of this capability. The same initial topic model is provided for all experiments.
Appendix B Experimental Settings and Interactive System Details
@@
B.1 Relational Approach Operational Details
@@
Finding Partitions Experts can find partitions in the space of unassigned instances. We currently support the K-means Jin2010 and Hierarchical Density-Based Clustering mendelsohn-etal-2021-modeling algorithms.
@@
Visualizing Global Explanations Experts can visualize aggregated statistics and explanations for the global state of the system. To do this, we aggregate all instances in the database. Explanations include theme distribution, coverage statistics, and t-sne plots JMLR:v9:vandermaaten08a.
@@
Table 3: Interactive Operations for the pacheco-2023-interactive System
Appendix C Semi-Structured Interview
C.1 Interviewing
We usually administered interviews after annotation sessions. In some cases, due to time constraints from the annotation session taking longer than expected, we were unable to pose these questions to the annotators. Asynchronous annotators were asked questions individually about their experience, whereas synchronous annotator groups were usually asked questions with their fellow annotators.
C.2 Script
@@
Appendix D LLM-based Experimental Configuration and Prompt Details
For this study, we use the same starting partitions as used in the relational approach experiment in App. B. The original work uses existing codes from a theoretical framework whereas we use codes defined by the user during this step, but this does not affect the overall process.
@@ Appendix D LLM-based Experimental Configuration and Prompt Details
Using the LLM, we generate 3 additional templates based on a prompt structure provided in the original work. We used the Llama 3.2 3B-Instruct (grattafiori2024llama3herdmodels) model for all generation tasks:
D.1 LLM Hyperparameters

### v5

@@ Appendix B Topic Model Experimental Settings
To ensure comparable results, we use the same starting distribution of 13 topics for all our experiments using the same hyperparameters as Fang et al. (2021), which are $\alpha=1.0$, $\beta=0.5$, $\gamma=1.5$. The QDTM also allows queries to be input prior to topic model initialization to produce partitions that follow prior knowledge, but we do not take advantage of this capability. The same initial topic model is provided for all experiments.
Appendix C Experimental Settings and Interactive System Details
@@
C.1 Relational Approach Operational Details
@@
Finding Partitions Experts can find partitions in the space of unassigned instances. We currently support the K-means Jin and Han (2010) and Hierarchical Density-Based Clustering Mendelsohn et al. (2021) algorithms.
@@
Visualizing Global Explanations Experts can visualize aggregated statistics and explanations for the global state of the system. To do this, we aggregate all instances in the database. Explanations include theme distribution, coverage statistics, and t-sne plots van der Maaten and Hinton (2008).
@@
Table 9: Interactive Operations for the Pacheco et al. (2023) System
Appendix D Semi-Structured Interview
D.1 Interviewing
We administered interviews after annotation sessions. Asynchronous annotators were asked questions individually about their experience, whereas synchronous annotator groups were asked questions with their fellow annotators.
D.2 Script
@@
Appendix E LLM-based Experimental Configuration and Prompt Details
For this study, we use the same starting partitions as used in the relational approach experiment in App. C. The original work uses existing codes from a theoretical framework whereas we use codes defined by the user during this step, but this does not affect the overall process.
@@ Appendix E LLM-based Experimental Configuration and Prompt Details
Using the LLM, we generate 3 additional templates based on a prompt structure provided in the original work. We used the Llama 3.2 3B-Instruct (Grattafiori et al., 2024) model for all generation tasks:
E.1 LLM Hyperparameters
@@ E.2 LLM Prompts
Appendix F Use of Pre-Existing Artifacts
All pre-existing artifacts utilized in this study, including datasets, software libraries, models, and computational tools, are publicly available under open-source and open-access licenses. This academic work adheres to all intended use guidelines and terms of service for the respective resources.